True or False Question

In order to establish a control chart for the mean of a process, 20 samples each of size 4 are collected. We note that ∑20 i=1 xi= 4000 and ∑20 i=1 si= 500. The value of the upper control limit of the chart for the mean is approximately equal to 204.5.

True False
Answer #1

Solution : False

Explanation :

Solution :- Given :- K=20 D=4 20 - 3 X = 4000 K=1 20 쿳 - 200 5 = 3 (si se) K ul 500 25 20 Control chart of Upper limit of mea

​​​​​​table values of control chart constant is;

X bar chart R chart S chart Subgroup A3 D3 B3 B. 3.267 2. 3 4 D 3.267 2.574 2.282 2.114 2.004 2.568 2.266 2.089 1.970 5 6 7 8